# Hydrodynamic memory and Quincke rotation

### v1

@@ Abstract
The spontaneous (so-called Quincke) rotation of an uncharged, solid, dielectric, spherical particle under a steady electric field is analyzed, accounting for the inertia of the particle and the transient fluid inertia, or “hydrodynamic memory,” due to the unsteady Stokes flow around the particle. The dynamics of the particle are encapsulated in three coupled nonlinear integro-differential equations for the evolution of the angular velocity of the particle, and the components of the induced dipole of the particle that are parallel and transverse to the applied field. These equations represent a generalization of the celebrated Lorenz system. A numerical solution of these ‘modified Lorenz equations’ (MLE) shows that hydrodynamic memory leads to an increase in the threshold field strength for chaotic particle rotation, which is in qualitative agreement with experimental observations. Furthermore, hydrodynamic memory leads to an increase in the range of field strengths where multi-stability between steady and chaotic rotation occurs. At large field strengths, chaos ceases and the particle is predicted to execute periodic rotational motion.
@@ I Introduction
The spontaneous rotary motion of a solid, dielectric particle in a low conductivity fluid under a direct current (DC) field was observed by Quincke in 1896 and is now commonly referred to as ‘Quincke rotation’ [1]. There has been a resurgence of interest in this phenomenon in recent years with applications in colloidal directed assembly [2, 3], droplet electro-hydrodynamics [4, 5, 6], active matter [7, 8, 9, 10, 11], and suspension rheology [12, 13, 14, 15] and conductivity [16]. For a sphere or long circular cylinder the rotation occurs only if the charge relaxation time of the particle, $\tau_{2}=\varepsilon_{2}/\sigma_{2}$, is greater than that of the fluid, $\tau_{1}=\varepsilon_{1}/\sigma_{1}$, where $\varepsilon$ is the permittivity and $\sigma$ is the electrical conductivity. (The subscripts 1 and 2 refer to the fluid and particle, respectively.) In this case, the base state has no fluid or particle motion and the induced dipole vector $\bm{p}$ inside the particle runs anti-parallel to applied field. Above a threshold field strength this mis-aligned dipole is unstable to infinitesimal perturbations and spontaneous rotation of the particle ensues, driven by an electric (Maxwell) torque arising from the induced dipole having a component transverse to the applied field. The angular velocity vector is orthogonal to the electric field; thus, a cylinder rotates about its axis of symmetry, and a sphere rotates about an arbitrary direction in the plane perpendicular to the field. In contrast, if $\tau_{2}<\tau_{1}$ the base state corresponds to an induced dipole parallel to the field, for which the electric torque dampens perturbations; this base state is stable.
@@
If the inertia of the particle and fluid are neglected, the angular momentum balance on the particle requires that the sum of electric and hydrodynamic torques is zero. In the absence of inertia, the hydrodynamic torque, $\bm{L}_{H}$, is given by the quasi-steady Stokes equations: for a sphere $\bm{L}_{H}=-8\pi\mu a^{3}\bm{\Omega}$, and for a cylinder the torque per unit length $\bm{L}_{H}=-4\pi\mu a^{2}\bm{\Omega}$, where $\mu$ is the fluid viscosity; $a$ is the radius for a sphere and cross-sectional radius for a cylinder; and $\bm{\Omega}$ is the angular velocity of the object. Consequently, for an inertialess sphere or cylinder in an inertialess fluid it is predicted that there is a single supercritical pitchfork bifurcation corresponding to the transition from a stationary base state to steady rotation. Note, the Qunicke rotation dynamics of a sphere and cylinder are mathematically equivalent, as will be shown later. In the absence of inertia, the critical field strengths for the onset of rotation for a long circular cylinder or sphere are the same and equal to $E_{c}=\sqrt{2\mu/[\varepsilon_{1}\tau_{MW}(\varepsilon_{21}-\sigma_{21})]}$, where for a sphere $\varepsilon_{21}=(\varepsilon_{2}-\varepsilon_{1})/(\varepsilon_{2}+2\varepsilon_{1})$ and $\sigma_{21}=(\sigma_{2}-\sigma_{1})/(\sigma_{2}+2\sigma_{1})$, and for a cylinder $\varepsilon_{21}=(\varepsilon_{2}-\varepsilon_{1})/(\varepsilon_{2}+\varepsilon_{1})$ and $\sigma_{21}=(\sigma_{2}-\sigma_{1})/(\sigma_{2}+\sigma_{1})$ [17]. Here, $\tau_{MW}$ is the Maxwell-Wagner relaxation time: for a sphere $\tau_{MW}=(\varepsilon_{2}+2\varepsilon_{1})/(\sigma_{2}+2\sigma_{1})$, and for a cylinder $\tau_{MW}=(\varepsilon_{2}+\varepsilon_{1})/(\sigma_{2}+\sigma_{1})$ [18]. Since $E_{c}$ is a real quantity this means rotation is possible only if $\varepsilon_{21}-\sigma_{21}>0$, which is equivalent to $\tau_{2}>\tau_{1}$. The rotation rate of the particle is $\Omega=\sqrt{[(E_{a}/E_{c})^{2}-1]}/\tau_{MW}$, where $E_{a}=|\bm{E}_{a}|$. Recall, the angular velocity is orthogonal to the applied field, $\bm{E}_{a}$, i.e. $\bm{E}_{a}\cdot\bm{\Omega}=0$. Therefore, the particle undergoes a supercritical pitchfork bifurcation at an applied field strength $E_{a}=E_{c}$: for $E_{a}<E_{c}$ there is a stable equilibrium in which the particle is non-rotating with an anti-parallel dipole; for $E_{a}>E_{c}$ this state is unstable and two stable equilibria correspond to a particle rotating at ±$\pm\Omega$ with a dipole that has a non-zero component transverse to the applied field. Which of the stable equilibria is attained, $+\Omega$ or $-\Omega$, is dictated by chance.
@@
Experiments by Lemaire and Lobry [19] and Peters et al. [20] observed, however, that the steady rotation is replaced by chaotic variations in the magnitude and direction of rotation above a second threshold field strength. Those authors showed that inclusion of particle inertia in the angular momentum balance leads to the Quincke rotation dynamics being described exactly by the celebrated Lorenz equations [21]. Inclusion of particle inertia does not affect the predicted value of $E_{c}$ or the steady rotation rate after the first, or primary, bifurcation. The Lorenz equations are well known to exhibit chaotic dynamics: in particular, they exhibit a secondary bifurcation from a non-trivial steady state to chaos, as observed in the Quincke rotor experiments [19, 20]. However, the measured threshold field strength for the secondary bifurcation (steady to chaotic rotation) is larger than that predicted from the Lorenz equations. Peters et al. [20] suggested that one source for this disagreement is their the neglect of fluid inertia. In experiments the Reynolds number associated with the particle rotation, $\mbox{{Re}}=a^{2}\Omega/\nu$, is small compared to unity. Here, $\Omega=|\bm{\Omega}|$ is the magnitude of the angular velocity and $\nu$ is the kinematic viscosity of the fluid. Consequently, the hydrodynamic torque is a linear function of the angular velocity. However, unsteady diffusion of fluid momentum arising from the angular acceleration of the particle could still be important. This means that the hydrodynamic torque is not an instantaneous function of angular velocity, as would be predicted from the quasi-steady Stokes equations. Instead, the unsteady Stokes equations must be used to calculate the hydrodynamic torque, which is now a function of the history of the angular velocity. Thereby, fluid inertia endows the particle with a hydrodynamic memory of its rotational motion.
@@
Hydrodynamic memory becomes important in Quincke rotation when the momentum diffusion time is comparable to the Maxwell-Wagner time for relaxation of the induced dipole. The momentum diffusion time is $\tau_{d}=a^{2}/\nu$. Thus, the dimensionless group $\gamma=\tau_{d}/\tau_{MW}$ characterizes the relevance of hydrodynamic memory: neglecting hydrodynamic memory is valid only if $\gamma\ll 1$; physically, here momentum diffuses quickly compared to the dipole relaxation. This condition is usually met for colloidal scale particles but not for larger, millimeter-sized, particles that are encountered in experiments and applications involving Quincke rotation [19, 20, 22, 23, 24]; here, hydrodynamic memory can be important since $\gamma$ is not small. As an example, Peters et al. [20] examined rotation of a glass capillary of $a=1\mathrm{mm}$ in transformer oil for which $\gamma=O(1)$; here, inertial effects result in chaotic rotation at sufficiently large fields. However, the mathematical model constructed by these authors included only the inertia of the particle, which is incomplete, as commented by themselves, for the following reason. Balancing the angular acceleration of the particle, which scales as $I\Omega/\tau_{p}$, where $I$ is the moment of inertia ($I=2ma^{2}/5$ for a sphere of mass $m$), with the quasi-steady hydrodynamic torque, which for a sphere scales as $8\pi\mu a^{3}\Omega$, gives the characteristic time for relaxation of the particle inertia as $\tau_{p}=(1/15)(\rho_{2}/\rho_{1})(a^{2}/\nu)$. This gives $\tau_{p}/\tau_{d}=(1/15)(\rho_{2}/\rho_{1})$, implying that the particle inertia and fluid inertia time scales are comparable, unless the particle is very much heavier or lighter than the fluid. Hence, one cannot accurately model inertial effects on Quincke rotation by only considering the inertia of the particle; fluid inertia via hydrodynamic memory must be accounted for. Therefore, our goal is to quantify the impact of hydrodynamic memory on Quincke rotation. To that end, we will derive a dynamical system for the rotor dynamics including hydrodynamic memory, in terms of the time evolution of the angular velocity of the particle and its dipole moment. The inclusion of hydrodynamic memory means that this dynamical system is integro-differential: we will call this set of equations as the ‘memory Lorenz equations’ (MLE), since they represent a generalization of the classic Lorenz system [21].
@@
In section II, we derive the MLE and then an ad hoc reduced form thereof that is amenable to analytical investigation, in which the memory kernel of the integro-differential term is simplified. In section III we conduct a linear stability analysis of the trivial solution of the MLE, showing that the first bifurcation is unaffected by hydrodynamic memory. In section IV we present a linear stability of the steady rotation, which indicates that hydrodynamic memory leads to an increase in the threshold field strength for which this solution becomes unstable. In section V, we develop a numerical scheme for integration of the MLE. In section VI, we explore the nonlinear dynamics beyond steady rotation and present results that show chaotic rotation is delayed until higher electric field strengths by hydrodynamic memory. We conclude in section VII.
@@ II Governing Equations
Here, we derive governing equations for the Quincke rotation of a solid spherical particle under a spatially uniform, DC electric field, accounting for hydrodynamic memory (figure 1). We use the Taylor-Melcher leaky dielectric model [25] in which charge transport is solely via Ohmic conduction in the bulk phases (particle and fluid), with charge localized at their interface. Let $\bm{r}$ denote the position vector from the centroid of the sphere and let $\phi(\bm{r})$ denote the electric potential. The leaky dielectric model stipulates that there are no free charges in the fluid and particle; hence, the electric potential satisfies Laplace’s equation, $\nabla^{2}\phi=0$. The applied field corresponds to the dipolar potential $-\bm{E}_{a}\!\cdot\!\bm{r}$. We define the disturbance potential in the fluid as $\phi_{1}(\bm{r})=\phi(\bm{r})+\bm{E}_{a}\!\cdot\!\bm{r}$ for $r=|\bm{r}|>a$ and the disturbance potential in the particle as $\phi_{2}(\bm{r})=\phi(\bm{r})+\bm{E}_{a}\!\cdot\!\bm{r}$ for $r=|\bm{r}|<a$. These disturbance potentials arise as the particle and fluid have different electrical properties, such that the applied field is distorted by the particle. The disturbance potentials satisfy $\nabla^{2}\phi_{1}=0$ and $\nabla^{2}\phi_{2}=0$, and $\phi_{1}$ decays as $r\to\infty$ and $\phi_{2}$ is bounded at $r=0$. The free charge is localized at the fluid-particle interface and represented by a surface charge distribution $q(\bm{r})$ at $r=a$. From Gauss’ law the surface charge is equated to the jump in dielectric displacement across the interface, $q(\bm{r})=-\bm{n}\!\cdot\!(\varepsilon_{1}\bm{\nabla}\phi_{1}-\varepsilon_{2}\bm{\nabla}\phi_{2})+\bm{n}\!\cdot\!\bm{E}_{a}(\varepsilon_{1}-\varepsilon_{2})$. Here, $\bm{n}$ is the outward unit vector to the particle surface. Additionally, the potential is continuous across the interface, $\phi_{1}(\bm{r})=\phi_{2}(\bm{r})$ at $r=a$. The final boundary condition enforces charge conservation at the interface,
@@
and $M(t)=0$ for $t<0$ (due to causality). Here $\mathrm{erfc}(z)=\textstyle{\frac{2}{\sqrt{\pi}}}\int_{z}^{\infty}\exp(-u^{2})\,\mathrm{d}u$ is the complimentary error function. Note that $M(t)=1/[3\sqrt{\pi t/\tau_{d}}]+O(1)$ as $t\to 0$ and $M(t)=(t/\tau_{d})^{-3/2}/6\sqrt{\pi}+O(t^{-5/2})$ as $t\to\infty$. Thus, the memory kernel is integrably singular at short times, which will require the use of special quadrature schemes. Together, (2) through (5) complete the formulation for Quincke rotation of a sphere with hydrodynamic memory. Evidently, the history of the angular acceleration must be accounted for, which means that the rotational dynamics obey a coupled set of nonlinear integro-differential equations.
To proceed, we adopt a Cartesian coordinate system anchored at the center of the particle, with the $z$ axis parallel to the applied field and the $x$ and $y$ axes in the plane perpendicular to the field. There are components of the dipole along each of these axes, say $p_{x}$, $p_{y}$, and $p_{z}$. From (3) we find $\bm{\Omega}\cdot\bm{E}_{a}=0$, so that the angular velocity only has non-zero components $\Omega_{x}$ and $\Omega_{y}$. Thus, it would seem that the rotational dynamics obey five coupled ordinary differential equations for the variables $p_{x}$, $p_{y}$, $p_{z}$, $\Omega_{x}$, and $\Omega_{y}$, as a function of time. However, these equations are invariant under the transformation $\{p_{x},p_{y},\Omega_{x},\Omega_{y}\}\to\{p_{y},p_{x},-\Omega_{y},-\Omega_{x}\}$, which suggests a solution with $p_{x}=p_{y}=p_{\perp}$ and $\Omega_{x}=-\Omega_{y}=-\Omega_{\perp}$. This means that the direction of rotation in the $x-y$ plane is arbitrary, and the strength of the transverse dipole, $p_{\perp}$, is independent of orientation in the $x-y$ plane. Thus, we have the dependent variables $p_{z},p_{\perp}$, and $\Omega_{\perp}$, which from (2) and (3) satisfy
@@
wherein the time derivatives are with respect to $s$; $M(s)=\textstyle{\frac{1}{3}}[1/\sqrt{\pi s}-\mathrm{erfc}(\sqrt{s/\gamma})\exp(s/\gamma)]$ is the normalized memory kernel; $Pr=8\pi\mu a^{3}\tau_{MW}/I$; and $r=(E_{a}/E_{c})^{2}$. Recall, $\gamma=\tau_{d}/\tau_{MW}$ is the ratio of momentum diffusion time to Maxwell-Wagner relaxation time. Hydrodynamic memory effects vanish as $\gamma\to 0$, in which case the integral term in (9) becomes negligible. In fact, without that integral term (9)-(11) are precisely the celebrated Lorenz equations [21]. The Lorenz equations also describe the Quincke rotation of a cylinder in an inertialess fluid: here, the normalized angular velocity $X$ is along the axis of the cylinder, and $Y$ and $Z$ are the transverse and parallel components of the dipole in the plane of the cross section of the cylinder [20]. Thus, the Qunicke rotation of a sphere and cylinder are mathematically equivalent. We refer to (9)-(11), with the integral term retained, as the ‘memory Lorenz equations’ (MLE). The MLE account for the inertia of the particle and surrounding fluid, both of which are important under physically realistic conditions.
II.1 Lorenz equations with simplified memory (sMLE)
To understand the role of memory in affecting the bifurcations of the Lorenz equations, we also consider simplified memory Lorenz equations (sMLE) with the memory kernel $M(s)=\alpha\,\text{e}^{-s/\alpha}$. This simplified memory kernel may be of practical relevance, in addition to mathematical convenience. Specifically, the popular Maxwell model stipulates the stress in a linear viscoelastic fluid to be a memory integral of the rate of strain, with a kernel that exponentially decays in time. Thus, the unsteady torque on a spherical particle executing time dependent rotation in a Maxwell fluid should have an exponentially fading memory of the history of its angular velocity, akin to (4). That is, the sMLE may be relevant to Quincke rotation in a linear viscoelastic fluid. Here, the memory is provided by viscoelasticity as opposed to fluid inertia.

### v2

@@ Abstract
The spontaneous (so-called Quincke) rotation of an uncharged, solid, dielectric, spherical particle under a steady uniform electric field is analyzed, accounting for the inertia of the particle and the transient fluid inertia, or “hydrodynamic memory,” due to the unsteady Stokes flow around the particle. The dynamics of the particle are encapsulated in three coupled nonlinear integro-differential equations for the evolution of the angular velocity of the particle, and the components of the induced dipole of the particle that are parallel and transverse to the applied field. These equations represent a generalization of the celebrated Lorenz system. A numerical solution of these ‘modified Lorenz equations’ (MLE) shows that hydrodynamic memory leads to an increase in the threshold field strength for chaotic particle rotation, which is in qualitative agreement with experimental observations. Furthermore, hydrodynamic memory leads to an increase in the range of field strengths where multi-stability between steady and chaotic rotation occurs. At large field strengths, chaos ceases and the particle is predicted to execute periodic rotational motion.
@@ I Introduction
The spontaneous rotary motion of a solid, dielectric particle in a low conductivity fluid under a direct current (DC) field was observed by Quincke in 1896 and is now commonly referred to as ‘Quincke rotation’ [1]. There has been a resurgence of interest in this phenomenon in recent years with applications in colloidal directed assembly [2, 3], droplet electro-hydrodynamics [4, 5, 6], and active matter [7, 8, 9, 10, 11]. Quincke rotation has also been shown to alter suspension rheology [12, 13, 14, 15] and conductivity [16]. For a sphere or long circular cylinder the rotation occurs only if the charge relaxation time of the particle, $\tau_{2}=\varepsilon_{2}/\sigma_{2}$, is greater than that of the fluid, $\tau_{1}=\varepsilon_{1}/\sigma_{1}$, where $\varepsilon$ is the permittivity and $\sigma$ is the electrical conductivity. Here, we consider that the field is applied perpendicular to the long axis of a cylinder, for which rotation occurs about that axis. (The subscripts 1 and 2 refer to the fluid and particle, respectively.) In this case, the base state has no fluid or particle motion and the induced dipole vector $\bm{p}$ inside the particle is oriented anti-parallel to applied field. Above a threshold field strength this mis-aligned dipole is unstable to infinitesimal perturbations and spontaneous rotation of the particle ensues, driven by an electric (Maxwell) torque arising from the induced dipole having a component transverse to the applied field. The angular velocity vector is orthogonal to the electric field; thus, a cylinder rotates about its axis of symmetry, and a sphere rotates about an arbitrary direction in the plane perpendicular to the field. In contrast, if $\tau_{2}<\tau_{1}$ the base state corresponds to an induced dipole parallel to the field, for which the electric torque dampens perturbations; this base state is stable.
@@
If the inertia of the particle and fluid are neglected, the angular momentum balance on the particle requires that the sum of electric and hydrodynamic torques is zero. In the absence of inertia, the hydrodynamic torque, $\bm{L}_{H}$, is given by the quasi-steady Stokes equations: for a sphere $\bm{L}_{H}=-8\pi\mu a^{3}\bm{\Omega}$, and for a cylinder the torque per unit length $\bm{L}_{H}=-4\pi\mu a^{2}\bm{\Omega}$, where $\mu$ is the fluid viscosity; $a$ is the radius for a sphere and cross-sectional radius for a cylinder; and $\bm{\Omega}$ is the angular velocity of the object. Consequently, for an inertialess sphere or cylinder in an inertialess fluid it is predicted that there is a single supercritical pitchfork bifurcation corresponding to the transition from a stationary base state to steady rotation. Note, the Quincke rotation dynamics of a sphere and cylinder are mathematically equivalent, as will be shown later. In the absence of inertia, the critical field strengths for the onset of rotation for a long circular cylinder or sphere have the same functional form of $E_{c}=\sqrt{2\mu/[\varepsilon_{1}\tau_{MW}(\varepsilon_{21}-\sigma_{21})]}$, where for a sphere $\varepsilon_{21}=(\varepsilon_{2}-\varepsilon_{1})/(\varepsilon_{2}+2\varepsilon_{1})$ and $\sigma_{21}=(\sigma_{2}-\sigma_{1})/(\sigma_{2}+2\sigma_{1})$, and for a cylinder $\varepsilon_{21}=(\varepsilon_{2}-\varepsilon_{1})/(\varepsilon_{2}+\varepsilon_{1})$ and $\sigma_{21}=(\sigma_{2}-\sigma_{1})/(\sigma_{2}+\sigma_{1})$ [17]. However, the numerical values for $E_{c}$ for a sphere and cylinder could be unequal due to the different definitions of $\varepsilon_{21}$ and $\sigma_{21}$ in each case. Here, $\tau_{MW}$ is the Maxwell-Wagner relaxation time: for a sphere $\tau_{MW}=(\varepsilon_{2}+2\varepsilon_{1})/(\sigma_{2}+2\sigma_{1})$, and for a cylinder $\tau_{MW}=(\varepsilon_{2}+\varepsilon_{1})/(\sigma_{2}+\sigma_{1})$ [18]. Since $E_{c}$ is a real quantity this means rotation is possible only if $\varepsilon_{21}-\sigma_{21}>0$, which is equivalent to $\tau_{2}>\tau_{1}$. The rotation rate of the particle is $\Omega=\sqrt{[(E_{a}/E_{c})^{2}-1]}/\tau_{MW}$, where $E_{a}=|\bm{E}_{a}|$. Recall, the angular velocity is orthogonal to the applied field, $\bm{E}_{a}$, i.e. $\bm{E}_{a}\cdot\bm{\Omega}=0$. Therefore, the particle undergoes a supercritical pitchfork bifurcation at an applied field strength $E_{a}=E_{c}$: for $E_{a}<E_{c}$ there is a stable equilibrium in which the particle is non-rotating with an anti-parallel dipole; for $E_{a}>E_{c}$ this state is unstable and two stable equilibria correspond to a particle rotating at $\pm\Omega$ with a dipole that has a non-zero component transverse to the applied field. Which of the stable equilibria is attained, $\Omega$ or $\Omega$, is dictated by chance.
@@
Experiments by Lemaire and Lobry [19] and Peters et al. [20] observed, however, that the steady rotation is replaced by chaotic variations in the magnitude and direction of rotation above a second threshold field strength. Those authors showed that inclusion of particle inertia in the angular momentum balance leads to the Quincke rotation dynamics being described exactly by the celebrated Lorenz equations [21]. Inclusion of particle inertia does not affect the predicted value of $E_{c}$ or the steady rotation rate after the first, or primary, bifurcation. The Lorenz equations are well known to exhibit chaotic dynamics: in particular, they exhibit a secondary bifurcation from a non-trivial steady state to chaos, as observed in the Quincke rotor experiments [19, 20]. However, the measured threshold field strength for the secondary bifurcation (steady to chaotic rotation) is larger than that predicted from the Lorenz equations. Peters et al. [20] suggested that one source of this disagreement is their neglect of fluid inertia. In experiments the Reynolds number associated with the particle rotation, $\mbox{{Re}}=a^{2}\Omega/\nu$, is small compared to unity. Here, $\Omega=|\bm{\Omega}|$ is the magnitude of the angular velocity and $\nu$ is the kinematic viscosity of the fluid. Consequently, the hydrodynamic torque is a linear function of the angular velocity. However, unsteady diffusion of fluid momentum arising from the angular acceleration of the particle could still be important. This means that the hydrodynamic torque is not an instantaneous function of angular velocity, as would be predicted from the quasi-steady Stokes equations. Instead, the unsteady Stokes equations must be used to calculate the hydrodynamic torque, which is now a function of the history of the angular velocity. Thereby, fluid inertia endows the particle with a hydrodynamic memory of its rotational motion.
@@
Hydrodynamic memory becomes important in Quincke rotation when the momentum diffusion time is comparable to the Maxwell-Wagner time for relaxation of the induced dipole. The momentum diffusion time is $\tau_{d}=a^{2}/\nu$. Thus, the dimensionless group $\gamma=\tau_{d}/\tau_{MW}$ characterizes the relevance of hydrodynamic memory: neglecting hydrodynamic memory is valid only if $\gamma\ll 1$; physically, here momentum diffuses quickly compared to the dipole relaxation. This condition is usually met for colloidal scale particles but not for larger, millimeter-sized, particles that are encountered in experiments and applications involving Quincke rotation [19, 20, 22, 23, 24]; here, hydrodynamic memory can be important since $\gamma$ is not small. As an example, Peters et al. [20] examined rotation of a glass capillary of $a=1\mathrm{mm}$ in transformer oil, for which they estimate $\tau_{d}=70\,\mathrm{ms}$ and $\tau_{MW}=150\,\mathrm{ms}$, yielding $\gamma=0.47$. Here, inertial effects result in chaotic rotation at sufficiently large fields. However, the mathematical model constructed by these authors included only the inertia of the particle, which is incomplete, as commented by themselves, for the following reason. Balancing the angular acceleration of the particle, which scales as $I\Omega/\tau_{p}$, where $I$ is the moment of inertia ($I=2ma^{2}/5$ for a sphere of mass $m$), with the quasi-steady hydrodynamic torque, which for a sphere scales as $8\pi\mu a^{3}\Omega$, gives the characteristic time for relaxation of the particle inertia as $\tau_{p}=(1/15)(\rho_{2}/\rho_{1})(a^{2}/\nu)$. This gives $\tau_{p}/\tau_{d}=(1/15)(\rho_{2}/\rho_{1})$, implying that the particle inertia and fluid inertia time scales are comparable, unless the particle is very much heavier or lighter than the fluid. Hence, one cannot accurately model inertial effects on Quincke rotation by only considering the inertia of the particle; fluid inertia via hydrodynamic memory must be accounted for. Therefore, our goal is to quantify the impact of hydrodynamic memory on Quincke rotation. To that end, we will derive a dynamical system for the rotor dynamics including hydrodynamic memory, in terms of the time evolution of the angular velocity of the particle and its dipole moment. The inclusion of hydrodynamic memory means that this dynamical system is integro-differential: we will call this set of equations as the ‘memory Lorenz equations’ (MLE), since they represent a generalization of the classic Lorenz system [21].
@@
In section II, we derive the MLE, and then in section III we derive an ad hoc reduced form thereof that is amenable to analytical investigation, in which the memory kernel of the integro-differential term is simplified. In section IV we conduct a linear stability analysis of the trivial solution of the MLE, showing that the first bifurcation is unaffected by hydrodynamic memory. In section V we present a linear stability of the steady rotation, which indicates that hydrodynamic memory leads to an increase in the threshold field strength for which this solution becomes unstable. In section VI, we develop a numerical scheme for integration of the MLE. In section VII, we explore the nonlinear dynamics beyond steady rotation and present results that show chaotic rotation is delayed until higher electric field strengths by hydrodynamic memory. We conclude in section VIII.
@@ II Governing Equations
Here, we derive governing equations for the Quincke rotation of a solid spherical particle under a spatially uniform, DC electric field, accounting for hydrodynamic memory (figure 1). We use the Taylor-Melcher leaky dielectric model [25] in which charge transport is solely via Ohmic conduction in the bulk phases (particle and fluid), with charge localized at their interface. Let $\bm{r}$ denote the position vector from the centroid of the sphere and let $\phi(\bm{r})$ denote the electric potential. The leaky dielectric model stipulates that there are no free charges in the fluid and particle; hence, the electric potential satisfies Laplace’s equation, $\nabla^{2}\phi=0$. The applied field corresponds to the dipolar potential $\bm{E}_{a}\!\cdot\!\bm{r}$. We define the disturbance potential in the fluid as $\phi_{1}(\bm{r})=\phi(\bm{r})+\bm{E}_{a}\!\cdot\!\bm{r}$ for $r=|\bm{r}|>a$ and the disturbance potential in the particle as $\phi_{2}(\bm{r})=\phi(\bm{r})+\bm{E}_{a}\!\cdot\!\bm{r}$ for $r=|\bm{r}|<a$. These disturbance potentials arise as the particle and fluid have different electrical properties, such that the applied field is distorted by the particle. The disturbance potentials satisfy $\nabla^{2}\phi_{1}=0$ and $\nabla^{2}\phi_{2}=0$, and $\phi_{1}$ decays as $r\to\infty$ and $\phi_{2}$ is bounded at $r=0$. The free charge is localized at the fluid-particle interface and represented by a surface charge distribution $q(\bm{r})$ at $r=a$. From Gauss’ law the surface charge is equated to the jump in dielectric displacement across the interface, $q(\bm{r})=-\bm{n}\!\cdot\!(\varepsilon_{1}\bm{\nabla}\phi_{1}-\varepsilon_{2}\bm{\nabla}\phi_{2})+\bm{n}\!\cdot\!\bm{E}_{a}(\varepsilon_{1}-\varepsilon_{2})$. Here, $\bm{n}$ is the outward unit vector to the particle surface. Additionally, the potential is continuous across the interface, $\phi_{1}(\bm{r})=\phi_{2}(\bm{r})$ at $r=a$. The final boundary condition enforces charge conservation at the interface,
@@
and $M(t)=0$ for $t<0$ (due to causality). Here $\mathrm{erfc}(z)=\textstyle{\frac{2}{\sqrt{\pi}}}\int_{z}^{\infty}\exp(-u^{2})\,\mathrm{d}u$ is the complimentary error function. Note that $M(t)=1/[3\sqrt{\pi t/\tau_{d}}]+O(1)$ as $t\to 0$ and $M(t)=(t/\tau_{d})^{-3/2}/(6\sqrt{\pi})+O(t^{-5/2})$ as $t\to\infty$. Thus, the memory kernel is integrably singular at short times, which will require the use of special quadrature schemes. Together, (2) through (5) complete the formulation for Quincke rotation of a sphere with hydrodynamic memory. Evidently, the history of the angular acceleration must be accounted for, which means that the rotational dynamics obey a coupled set of nonlinear integro-differential equations.
To proceed, we adopt a Cartesian coordinate system anchored at the center of the particle, with the $z$ axis parallel to the applied field and the $x$ and $y$ axes in the plane perpendicular to the field. There are components of the dipole along each of these axes, say $p_{x}$, $p_{y}$, and $p_{z}$. From (3) we find $\bm{\Omega}\cdot\bm{E}_{a}=0$, so that the angular velocity only has non-zero components $\Omega_{x}$ and $\Omega_{y}$. Thus, it would seem that the rotational dynamics obey five coupled ordinary differential equations for the variables $p_{x}$, $p_{y}$, $p_{z}$, $\Omega_{x}$, and $\Omega_{y}$, as a function of time. However, these equations are invariant under the transformation $\{p_{x},p_{y},\Omega_{x},\Omega_{y}\}\to\{p_{y},p_{x},-\Omega_{y},-\Omega_{x}\}$, which suggests a solution with $p_{x}=p_{y}=p_{\perp}$ and $\Omega_{x}=-\Omega_{y}=-\Omega_{\perp}$. This means that the direction of rotation in the xyx-y plane is arbitrary, and the strength of the transverse dipole, $p_{\perp}$, is independent of orientation in the xyx-y plane. Thus, we have the dependent variables $p_{z},p_{\perp}$, and $\Omega_{\perp}$, which from (2) and (3) satisfy
@@
wherein the time derivatives are with respect to $s$; $M(s)=\textstyle{\frac{1}{3}}[1/\sqrt{\pi s}-\mathrm{erfc}(\sqrt{s/\gamma})\exp(s/\gamma)]$ is the normalized memory kernel; $Pr=8\pi\mu a^{3}\tau_{MW}/I$; and $r=(E_{a}/E_{c})^{2}$. Recall, $\gamma=\tau_{d}/\tau_{MW}$ is the ratio of momentum diffusion time to Maxwell-Wagner relaxation time. Hydrodynamic memory effects vanish as $\gamma\to 0$, in which case the integral term in (9) becomes negligible. In fact, without that integral term (9)-(11) are precisely the celebrated Lorenz equations [21]. The Lorenz equations also describe the Quincke rotation of a cylinder in an inertialess fluid: here, the normalized angular velocity $X$ is along the axis of the cylinder, and $Y$ and $Z$ are the transverse and parallel components of the dipole in the plane of the cross section of the cylinder [20]. Thus, the Quincke rotation of a sphere and cylinder are mathematically equivalent. We refer to (9)-(11), with the integral term retained, as the ‘memory Lorenz equations’ (MLE). The MLE account for the inertia of the particle and surrounding fluid, both of which are important under physically realistic conditions.
III Lorenz equations with simplified memory (sMLE)
To understand more generally the role of memory in affecting the bifurcations of the Lorenz equations, we introduce simplified memory Lorenz equations (sMLE) with the memory kernel $M(s)=\alpha\,\text{e}^{-s/\alpha}$. This simplified memory kernel may be of practical relevance, in addition to mathematical convenience. Specifically, the stress in a linear viscoelastic fluid is a memory integral of the rate of strain, which captures the physical effect that viscoelasticity, akin to inertia, endows a fluid with a memory of the history of its deformation. For the popular Maxwell model the kernel in the memory integral decays exponentially in time. Thus, the unsteady torque on a spherical particle executing time dependent rotation in a Maxwell fluid should have an exponentially fading memory of the history of its angular velocity, akin to (4). That is, the sMLE may be relevant to Quincke rotation in a linear viscoelastic fluid. Here, the memory is provided by viscoelasticity as opposed to fluid inertia. The dimensionless parameter $\alpha$ in the sMLE would physically correspond to the ratio of the viscoelastic relaxation time of the fluid to the Maxwell-Wagner relaxation time. That is, $\alpha$ plays an analogous role to $\gamma$ in the MLE.